# Lattice Random Walk Discretisations of Stochastic Differential Equations

### v1

@@ II.2 Selecting $\delta_{x}$
This feasibility condition makes the stochastic nature of the discretisation apparent. Since we need $\delta_{t}>0$, (8) implies we also need $\sigma(x,t)^{2}>0$, thus the discretisation is inherently stochastic and does not reduce to an ODE discretisation in the same way Euler-Maruyama does with $\sigma(x,t)=0$.
@@
For drift or diffusion functions that have a large dynamic range over the course of a SDE trajectory, we can also generalize $\delta_{x}\to\delta_{x}(t,x)$ to be a function of $t$ or even $x$. This is particularly useful for SDEs with time-varying diffusion functions, such as variance-exploding diffusion models [51].
@@ IV.1 Stochastic computing
As an example of the potential efficiencies of stochastic computing, consider a stochastic (unipolar) number $x$ (that is a random bit generated from $\text{Bernoulli}(x)$) and a stochastic number $y$ then a new stochastic number representing the product $xy$ can be computed with a simple AND gate
$\text{AND}(x,y)=xy.$
Similarly we have $\text{NAND}(x,y)=1-xy$ and other basic gates, see Chapter 5 in [16] (although the NAND gate is universal for boolean logic). This makes linear operations (such as matrix multiplications) in stochastic computing extremely efficient and can be applied in a single input bit to single output bit model of computation. And importantly avoid any floating point arithmetic or quantisation error beyond that of the original stochastic numbers generation.
@@ V.2 Poisson random effects model (non-globally Lipshitz)
Figure 5: Robustness to stepsize for a non-Lipschitz Poisson random effects model. Error from the ergodic mean of the generated samples and the true parameter, averaged over 100 different seeds.
@@ V.3 Diffusion model
In this experiment we deploy lattice random walk discretisation to a large-scale state-of-the-art image generation model. We use the popular Stable Diffusion 3.5 model [12] which has over 8 billion parameters. Stable Diffusion 3.5 is technically a flow-matching model [27] and differs in training than that of a continuous-time diffusion model [51]. This class of models are usually viewed as a deterministic ordinary differential equation at inference time. However, as described in Appendix D, flow-matching models can be recast as SDEs of the form
@@
In Figure 6 we compare the Euler-Maruyama and LRW discretisations for generating prompted images from the pretrained Stable Diffusion 3.5 model [12], which provide $s(x,t)$ and $\varsigma(t)$ in (13). In all cases, we set the level of noise $\alpha(t)=a\dot{\varsigma}(t)\varsigma(t)$ following [27, 51] with $a=0.3$. In this case, unlike the previous experiments, the diffusion term $\sigma(t)=\sqrt{2\alpha(t)}$ is time-varying, but we can still apply the rule of thumb from Section II.2 with a time-varying $\delta_{x}(t)=\sqrt{\delta_{t}}\sigma(t)$.
@@ VI.2 Future work
Additionally, the development of specialized stochastic computing hardware implementations represents perhaps the most transformative direction and one we are actively pursuing [41, 4, 9, 1]. The discrete nature of LRW outputs makes it ideally suited for stochastic computing architectures. Indeed, developing dedicated hardware that can efficiently generate and manipulate the binary/ternary random variables could unlock many orders of magnitude speedup for SDE simulation in applications such as large-scale diffusion models, as unlocked by the introduction of the lattice random walk discretisation.
@@ Acknowledgments
We thank Lars Holdijk, Rob Brekelmans, Gavin Crooks, and Max Welling for valuable feedback during the writing of the manuscript. We thank the Advanced Research and Invention Agency’s (ARIA) Scaling Compute programme for funding this work.
@@ Appendix A Weak convergence
Let $\varphi\colon\mathbb{R}^{d}\to\mathbb{R}$ be a sufficiently smooth test function and denote by $X(t)$ the true solution of (1) with diagonal $\sigma(x,t)$. Then the generator of the true solution has the form [46, page 50]
@@
From the one-step expansion we have
@@
i.e. the scheme is weak-order 1 for $\delta_{x,i}=\Theta(\delta_{t}^{1/2})$.
@@ Appendix D From Flows to SDEs
The output of the flow matching training is a velocity $u_{\theta}(x_{\tau},\tau)$ that can be used within an ordinary differential equation (ODE) solver that reverses the process for $t=0\to 1$
@@
As described in [27], the trained flow matching model actually recovers a scaled version of the score function $s(x,t)=\nabla_{x}\log p_{t}(x,t)$. That is, the ordinary differential equation, now moving forward in time $t=0\to 1$, can be written as
$dx=\dot{\varsigma}(t)\varsigma(t)\nabla_{x}\log p_{t}(x_{t},t)dt.$
and we can convert the velocity to the score with knowledge of the noise schedule $\varsigma(t)$
$\nabla_{x}\log p_{t}(x_{t},t)=-\frac{1}{\dot{\varsigma}(t)\varsigma(t)}u_{\theta}(x_{t},1-t).$
Further this ODE can be generalised to an SDE by adding any amount of Langevin noise that is invariant for the instantenous distribution $p(x,t)$ [39]
$dx_{t}=\dot{\varsigma}(t)\varsigma(t)\nabla_{x}\log p_{t}(x_{t},t)dt+\alpha(t)\nabla_{x}\log p_{t}(x_{t},t)dt+\sqrt{2\alpha(t)}dw_{t},$
matching that of (13) where $\alpha(t)\geq 0$ controls the level of added noise but remarkably does not affect the marginal distributions $p(x,t)$ [27].

### v2

@@ I Introduction
The LRW discretization as well as Euler-Maruyama and a continuous-time trajectory are visualized on a toy two-dimensional distribution in Fig. 1, where one can clearly see the drastic difference in approaches between assuming infinite precision and traversing a lattice.
@@ II.2 Selecting $\delta_{x}$
This feasibility condition makes the stochastic nature of the discretisation apparent. Since we need $\delta_{t}>0$, (8) implies we also need $\sigma(x,t)^{2}>0$, thus the discretisation is inherently stochastic and does not reduce to an ordinary differential equation (ODE) discretisation in the same way Euler-Maruyama does with $\sigma(x,t)=0$. However, as we will see in Section II.4, we can adjust the LRW discretisation into one that is consistent for deterministic ODEs.
@@
For drift or diffusion functions that have a large dynamic range over the course of a SDE trajectory, we can also generalize $\delta_{x}\to\delta_{x}(t,x)$ to be a function of $t$ or even $x$ so long as we still have $\delta_{x,i}(x,t)=\Theta(\sqrt{\delta_{t}})$. This is particularly useful for SDEs with time-varying diffusion functions, such as variance-exploding diffusion models [56].
@@ II.3 Diagonal Diffusion Assumption
II.4 Lattice Random Walk for Ordinary Differential Equations
An alternative LRW discretisation that is consistent for ODEs can be achieved with a variance that is quadratic in the stepsize $O(\delta_{t}^{2})$ since then the variance decays faster than the mean as $\delta_{t}\to 0$.
Specifically if the increment’s moments are modified to have the form
$\begin{split}\mathbb{E}[x_{t+\delta_{t}}-x_{t}\mid x_{t}]&=\delta_{t}f(x_{t},t),\\ \mathbb{E}[(x_{t+\delta_{t}}-x_{t})^{2}\mid x_{t}]&=\delta_{t}^{2}g(x_{t},t)^{2}.\end{split}$ (10)
Then for some $g(x,t)$ independent of $\delta_{t}$ the discretisation will converge to the ODE $dx=f(x,t)dt$ as $\delta_{t}\to 0$.
Since ODEs are deterministic, we don’t typically differentiate between strong and weak error, since they coincide. In this case, we don’t necessarily expect the above LRW discretisation to have first order error $O(\delta_{t})$ because of the required injected noise. However, for many applications, such as Bayesian sampling and inference in diffusion models in Section V.3 a version of weak error may still apply for ODEs where expectations are taken with respect to a random initial distribution $p(x_{0})$. A full error analysis in this setting of ODE discretisation is left for future work.
@@ IV.1 Stochastic Computing
As an example of the potential efficiencies of stochastic computing, consider a stochastic (unipolar) number $x$ (that is a random bit generated from $\text{Bernoulli}(x)$) and a stochastic number $y$ then a new stochastic number representing the product $xy$ can be computed with a simple AND gate $\text{AND}(x,y)=xy$. Similarly we have $\text{NAND}(x,y)=1-xy$ and other basic gates, see Chapter 5 in [21] (although the NAND gate is universal for boolean logic). This makes linear operations (such as matrix multiplications) in stochastic computing extremely efficient and can be applied in a single input bit to single output bit model of computation. And importantly avoid any floating point arithmetic or quantisation error beyond that of the original stochastic numbers generation.
@@
In Appendix F, we give a concrete stochastic computing protocol for LRW on linear SDEs, i.e. Ornstein-Uhlenbeck (OU) processes $dx=-(Ax-b)dt+\sigma dw$. It utilises a variant on the commonly used stochastic computing multiplexer (MUX) primitive for scaled addition [21]. By computing alias tables [57] in a fast preprocessing step, categorical sampling can be performed in $O(1)$ time, enabling a single iteration of LRW with work $O(d)$ and parallel time $O(1)$ versus traditional matrix-vector product arithmetic which has work $O(d^{2})$ and parallel time $O(\log d)$ [5]. The overall complexities required to achieve weak error $\epsilon$ are detailed in Table 2, which in both cases requires $O(\epsilon^{-1})$ steps since both EM and LRW have weak order 1.
We note several caveats to this protocol: (i) the MUX-based approach imposes a maximal stepsize constraint (see Appendix F), (ii) the protocol is specific to linear OU processes and does not directly extend to non-linear drifts, and (iii) OU processes can alternatively be sampled exactly with $O(d^{3})$ preprocessing via matrix decomposition [15], though this requires significantly more complex floating-point operations whilst higher-order discretisations (see Table 1) can also improve the dependency on $\epsilon$.
The key advantage of the stochastic computing approach is that it replaces floating-point matrix-vector products with simple bit-level operations, potentially enabling significant energy and latency improvements on specialized hardware. Notably CPUs and GPUs are designed for dense floating point arithmetic whereas the stochastic approach moves the bottleneck to repeated categorical sampling via alias tables which represents a fundamentally different pipeline ripe for specialized hardware acceleration.
Work
Parallel
Span Time
Required
Threads
Floating Point EM $O(\epsilon^{-1}d^{2})$ $O(\epsilon^{-1}\log d)$ $O(d^{2})$
Stochastic LRW $O(d^{2}+\epsilon^{-1}d)$ $O(d+\epsilon^{-1})$ $O(d)$
Table 2: Complexity of sampling from an OU process with traditional vs stochastic computing protocols. We differentiate between work (the total number of operations), parallel span time (the time taken to perform the operations assuming sufficient parallelism) and required threads (the number of threads required to achieve the stated parallel span). The table shows the complexities required to achieve weak error $\epsilon$ on a dense OU process with $d$ dimensions, including both preprocessing and iterative computation.
@@ V.2 Poisson Random Effects Model (Non-globally Lipshitz)
Figure 5: Robustness to stepsize for a non-Lipschitz Poisson random effects model. Error from the ergodic mean of the generated samples and the true parameter, averaged over 50 different seeds.
@@ V.3 Diffusion Model
In this experiment we deploy lattice random walk discretisation to a large-scale state-of-the-art image generation model. We use the popular Stable Diffusion 3.5 model [17] which has over 8 billion parameters. Stable Diffusion 3.5 is technically a flow-matching model [33] and differs in training than that of a continuous-time diffusion model [56]. This class of models are usually viewed as a deterministic ordinary differential equation at inference time. However, as described in Appendix D (which expands on the exposition from [33]), flow-matching models can be recast as SDEs of the form
@@
In Figure 6 we compare the Euler-Maruyama and LRW discretisations for generating prompted images from the pretrained Stable Diffusion 3.5 model [17], which provide $s(x,t)$ and $\varsigma(t)$ in (14). In all cases, we set the level of noise $\alpha(t)=-a\dot{\varsigma}(t)\varsigma(t)$ following [33, 56] with $a=0.3$. In this case, unlike the previous experiments, the diffusion term $\sigma(t)=\sqrt{2\alpha(t)}$ is time-varying, but we can still apply the rule of thumb from Section II.2 with a time-varying $\delta_{x}(t)=\sqrt{\delta_{t}}\sigma(t)$.
A quantitative numerical analysis comparing Fréchet inception distance between EM and LRW over varying steps is found in Appendix E.
@@ VI.2 Future Work
Additionally, the development of specialized stochastic computing hardware implementations represents perhaps the most transformative direction and one we are actively pursuing [46, 4, 10, 1, 13, 12]. The discrete nature of LRW outputs makes it ideally suited for stochastic computing architectures. Indeed, developing dedicated hardware that can efficiently generate and manipulate the binary/ternary random variables could unlock many orders of magnitude speedup for SDE simulation in applications such as large-scale diffusion models, as unlocked by the introduction of the lattice random walk discretisation.
@@ Acknowledgments
We thank Lars Holdijk, Rob Brekelmans, Gavin Crooks, Jan Ole Ernst, and Max Welling for valuable feedback during the writing of the manuscript. We thank the Advanced Research and Invention Agency’s (ARIA) Scaling Compute programme for funding this work.
Author contributions
S.D. invented the binary form of the algorithm. M.A. derived the binary discretization for unit step size. M.A. assisted with the derivation to the more general ternary form. S.D. and D.M. ran the experiments. M.A., Z.B. and P.C. contributed to fundamental early work on a version for linear systems and its connection to stochastic computing. All authors reviewed the manuscript.
Competing Interests
All authors declare no financial or non-financial competing interests.
@@ Appendix A Weak Convergence
Let $\varphi\colon\mathbb{R}^{d}\to\mathbb{R}$ be a sufficiently smooth test function and denote by $X(t)$ the true solution of (1) with diagonal $\sigma(x,t)$. Then the generator of the true solution has the form [51, page 50]
@@
From the one‐step expansion we have
@@
i.e. the scheme is weak‐order 1 for $\delta_{x,i}=\Theta(\delta_{t}^{1/2})$.
@@ Appendix D From Flows to SDEs
In this section we summarize the framework introduced in [33] unifying diffusion and flow models at inference time as a family of SDEs with flexible amount of noise (including zero noise i.e. an ODE as a special case).
@@
The output of the flow matching training is a velocity $u_{\theta}(x_{\tau},\tau)$ that can be used within an ordinary differential equation (ODE) solver that reverses the process for $\tau=1\to 0$
@@
As described in [33], the trained flow matching model actually recovers a scaled version of the score function $s(x,t)=\nabla_{x}\log p_{t}(x,t)$. That is, the ordinary differential equation, moving forward in time $t=0\to 1$, can be written as
$dx=-\dot{\varsigma}(t)\varsigma(t)\nabla_{x}\log p_{t}(x_{t},t)dt,$
where $\varsigma(t)=\sigma(1-t)$. Since $\sigma(\tau)$ is typically increasing with $\tau$ then $\varsigma(t)$ is decreasing with $t$ and $\dot{\varsigma}(t)$ is negative.
We can convert the velocity to the score with knowledge of the noise schedule $\varsigma(t)$
$\nabla_{x}\log p_{t}(x_{t},t)=\frac{1}{\dot{\varsigma}(t)\varsigma(t)}u_{\theta}(x_{t},1-t).$
Further, this ODE can be generalised to an SDE by adding any amount of Langevin noise that is invariant for the instantaneous distribution $p(x,t)$ [44]
$dx_{t}=-\dot{\varsigma}(t)\varsigma(t)\nabla_{x}\log p_{t}(x_{t},t)dt+\alpha(t)\nabla_{x}\log p_{t}(x_{t},t)dt+\sqrt{2\alpha(t)}dw_{t},$
matching that of (14) where $\alpha(t)\geq 0$ controls the level of added noise but remarkably does not affect the marginal distributions $p(x,t)$ [33, 16]. A common choice is $\alpha(t)=-a\dot{\varsigma}(t)\varsigma(t)$ for some value of $a>0$ (noting that $\dot{\varsigma}(t)<0$ implies $\alpha(t)>0$).
Appendix E Fréchet Inception Distance Experiment
In Figure 7, we extend the experiment in Section V.3 to give a numerical quantification of the quality of images generated by LRW - we use the commonly used Fréchet Inception distance to numerically quantify images generated with gold standard images (in this case running SD3.5 with a small stepsize over 100 steps). In each case we generate 100 images using the same prompt as in Section V.3. We see that Euler-Maruyama out-performs at low number of steps (and therefore large stepsize), with LRW catching up for more steps and converging at a similar rate. This behaviour is not unsurprising given the harsh quantisation of the LRW increment and indeed it is somewhat remarkable that LRW performance is close - noting that after 50 steps EM and LRW are not visually distinguishable as seen in Figure 6a.
Figure 7: FID distance from gold-standard images All images were generated with Stable Diffusion 3.5 with the prompt “A kitten riding a skateboard holding a cup of tea” as in Figure 6.
Appendix F Simlulating Ornstein-Uhlenbeck Processes with Stochastic Computing
As a concrete example of the LRW discretisation with stochastic computing, we can consider simulating an Ornstein-Uhlenbeck process
$dx=-(Ax-b)dt+\sigma dw.$
An iteration of binary LRW amounts to sampling a bipolar vector $v_{t}\in\{-1,1\}^{d}$ with moments
$\mathbb{E}[v_{t}]=-\sigma^{-1}\sqrt{\delta_{t}}(Ax_{t}-b),\quad\mathbb{E}[v_{t}^{2}]=1,$
then we get LRW update $x_{t+\delta_{t}}=x_{t}+\sqrt{\delta_{t}}\sigma v_{t}$.
The bipolar increment expectation can be written as a single matrix vector product
$\mathbb{E}[v_{t}]=-\sigma^{-1}\sqrt{\delta_{t}}By_{t},\quad B=(A\,,\;{-}b),\quad y_{t}=(x_{t}^{\top}\,,\;1)^{\top}.$ (18)
A common method for (scaled) addition in stochastic computing is the so-called multiplexer approach (e.g. Section 3 in [21]). Generalizing to a dot product we can write the protocol for input stochastic bipolar vector $Y\in\{-1,1\}^{n}$ with expected value $\mathbb{E}[Y_{i}]=y_{i}$ and real weights $w\in\mathbb{R}^{n}$ as
$k$ $\sim\text{Categorical}\left(\frac{|w_{1}|}{\|w\|_{1}},\dots,\frac{|w_{n}|}{\|w\|_{1}}\right),$
$b$ $=\text{sign}(w_{k})Y_{k},$
where $\|w\|_{1}=\sum_{j}|w_{j}|$. This gives expected value as the (scaled) dot product
$\mathbb{E}[b]=\frac{w^{\top}y}{\|w\|_{1}},$
which can be repeated in parallel for the $d$ rows of the matrix-vector product in (18).
To apply this protocol, we encode $y_{t}$ as a stochastic bipolar vector: assuming $\|y_{t}\|_{\infty}\leq M$, each component $y_{t,j}$ is represented by a random bit $Y_{j}\in\{-1,+1\}$ with $\mathbb{P}[Y_{j}=+1]=(1+y_{t,j}/M)/2$, giving $\mathbb{E}[Y_{j}]=y_{t,j}/M$. Sampling from the categorical distribution naïvely requires $O(n)$ time, but using the alias method [57] with $O(n)$ preprocessing, each sample can be drawn in $O(1)$ time. This gives $O(d^{2})$ preprocessing (for $d$ alias tables, one per row of $B$) and $O(d)$ work per LRW iteration, or $O(1)$ parallel time with $O(d)$ threads since the $d$ categorical samples are independent.
For the protocol to be valid, we require $|\mathbb{E}[v_{t,i}]|\leq 1$ for each coordinate $i$. Combined with a bound $\|y_{t}\|_{\infty}\leq M$ (which holds if the OU process remains bounded), we obtain a maximal stepsize
$\delta_{t}$ $\leq\left(\frac{\sigma}{M\bar{B}}\right)^{2},$
where $\bar{B}=\max_{i}\sum_{j}|B_{ij}|$ is the maximum absolute row sum of $B$.